Father & son present age 60. 10 year after father age son age 3 malti. Father & son present age?
Answer by ewatrrr(24785)

NOW
son x and Father (60-x)
In Ten years...
(60-x) + 10 = 3(x+10)
70 - x = 3x + 30
40 = 2x
20 = x, son's age NOW and Father is 40 60-20